# Zhou, Orfanoudaki, and Zhu

### v2

@@
In many operational settings, decision-makers must commit to actions before uncertainty resolves, but existing optimization tools rarely quantify how consistently a chosen decision remains optimal across plausible futures. This paper introduces CREDO—Conformalized Risk Estimation for Decision Optimization, a distribution-free framework that quantifies the probability that a prescribed decision remains (near-)optimal across realizations of uncertainty. CREDO reformulates decision risk through the inverse feasible region—the set of outcomes under which a decision is optimal—and estimates its probability using inner approximations constructed from conformal prediction balls generated by a conditional generative model. By calibrating each ball to lie entirely within the inverse feasible region, CREDO obtains finite-sample valid lower bounds on decision optimality without parametric assumptions. The method avoids the conservatism of worst-case robust optimization, is compatible with modern generative models, and applies broadly to convex optimization problems. We establish theoretical validity guarantees, develop efficient computational procedures, and demonstrate through extensive numerical experiments that CREDO provides accurate, interpretable, and reliable assessments of decision reliability in both synthetic and application-motivated settings.
@@ 1 Introduction
To operationalize this perspective, we introduce CREDO—Conformalized Risk Estimation for Decision Optimization, a framework that quantifies, for any candidate decision, a distribution-free lower bound on the probability that it remains (near-)optimal (see Figure 1 for an illustration). The proposed approach integrates inverse optimization (chan2025inverse) with distribution-free uncertainty quantification via conformal prediction (CP) (angelopoulos2021gentle). The core challenge is to evaluate the likelihood that an uncertain scenario falls within the decision’s inverse feasible region, namely the set of scenarios under which that decision remains optimal or near-optimal, and directly characterizing this region is generally intractable due to its implicit and complex geometry.
@@
Illustration of Conformalized Decision Risk Assessment (CREDO) compared against Prediction-Then-Optimize (PTO). For a given candidate decision $z$ and dataset $\{(x_{i},y_{i})\}_{i=1}^{n}$, CREDO aims to estimate the decision risk, which is defined as the probability of $z$ to be suboptimal. CREDO finds the inverse feasible region using inverse optimization $\pi^{-1}$, and uses a generative model to simulate multiple scenarios, which are then calibrated to produce the conformalized risk estimate $\hat{\alpha}$.
CREDO circumvents this challenge by constructing inner approximations of the inverse feasible region using multiple CP balls generated from a conditional generative model. Each ball serves as a finite-sample valid prediction region for the outcome. By shrinking the radius until the ball fits entirely inside the inverse feasible region, we obtain a certified lower bound on the decision’s optimality probability. Aggregating these calibrated bounds yields a distribution-free, finite-sample valid estimator of decision risk. CREDO is compatible with modern generative forecasting models, supports multiple conformal radius constructions, and applies broadly to convex decision problems. We establish theoretical validity for the estimator, develop efficient computational procedures, and demonstrate through extensive numerical experiments that CREDO provides accurate, interpretable, and reliable assessments of decision reliability across both synthetic and application-driven settings. In what follows, we summarize the key contributions of our work:
  • Problem Formulation: We formulate the decision risk assessment as estimating the probability that a candidate decision remains (near-)optimal under uncertainty. By expressing this probability through inverse feasible regions, we convert an otherwise intractable evaluation problem into a tractable estimation task, revealing a fundamental link between inverse optimization and distribution-free uncertainty quantification.
  • CREDO Framework: We introduce CREDO, a distribution-free framework based on CP that delivers finite-sample valid lower bounds on decision optimality, and establish marginal conservativeness, asymptotic consistency, and accuracy guarantees.
@@
  • Empirical Validation: Through controlled synthetic experiments and a real-world power grid planning application with 1,700+ solar installations, we verify CREDO’s theoretical guarantees and demonstrate its practical advantages. Our method achieves 100% empirical validity under conservative radius choices, superior true positive rates (up to 78.75% improvement over baselines), and consistently selects decisions with higher empirical confidence rankings than existing prescriptive methods.
The remainder of this paper is organized as follows. Section 2 reviews the related literature and positions our contributions. Section 3 formalizes the decision risk assessment problem and provides motivating examples. In Section 4, we present the CREDO methodology, combining inverse optimization with generative CP. Section 5 establishes theoretical guarantees including validity, consistency, and true positive rate analysis. Section 6 details the proposed computational implementation of CREDO, while Section 7 demonstrates its effectiveness through synthetic and real-world experiments. Section 8 concludes with discussions of implications and future work. Technical proofs and additional details are in the Electronic Supplement.
@@ 2 Related Work
Our work draws from and contributes to four distinct research streams: ($i$) decision-making under uncertainty in operations research, ($ii$) distribution-free uncertainty quantification through CP, (iiiiii) inverse optimization for understanding decision structures, and ($iv$) human-algorithm collaboration in decision support systems. In what follows, we position the CREDO approach relative to these lines of research.
@@
Classical approaches to optimization under uncertainty can be categorized by their treatment of unknown parameters. Stochastic optimization seeks decisions that minimize the expected value of the objective function across probabilistic scenarios, typically implemented through sample average approximation over historical data (shapiro2021lectures, kleywegt2002sample, lan2025bias). The predict-then-optimize paradigm offers an alternative by first estimating unknown parameters via predictive models, then solving the resulting deterministic optimization problem (bertsimas2020predictive, elmachtoub2022smart). Acknowledging that probability distributions themselves may be misspecified, robust optimization is employed to hedge against worst-case realizations within specified uncertainty sets (bertsimas2006robust, ben2002robust). Distributionally robust optimization generalizes this concept further, considering worst-case probability distributions rather than point realizations, thereby balancing conservatism with statistical plausibility (delage2010distributionally, rahimian2022frameworks). Recognizing that prediction errors can compound when models are trained separately from their downstream use, decision-focused learning trains predictive models end-to-end by differentiating through the optimization layer, directly minimizing downstream decision costs rather than prediction errors (amos2017optnet, mandi2024decision). While these prescriptive paradigms effectively recommend decisions, our work adopts a parallel goal: rather than prescribing a decision, we focus on auditing decisions by providing a quantitative, data-driven assessment of the risk associated with taking a given decision. This perspective complements existing decision-prescription methods and even human judgment by adding an additional layer of transparency and accountability, thereby supporting more reliable and defensible decision-making.
@@
Methodologically, our work builds upon the CP literature, a distribution-free approach for uncertainty quantification through calibrated prediction sets (shafer2008tutorial, vovk2005algorithmic). We highlight two strands of recent research that are particularly relevant to our work. First, generative conformal methods draw multiple samples from generative models to construct both the calibration sets and the resulting CP regions, thus significantly improving efficiency (i.e., tightness of the prediction sets) especially when the data distribution is highly complex or dispersed (zheng2024optimizing, zhou2024hierarchical). Second, inverse CP estimates miscoverage rates for fixed prediction sets by identifying the smallest miscoverage level at which conformal regions are contained within the target set (prinster2023jaws, singh2024distribution). Advances in $e$-value CP strengthen this approach by enabling post-hoc selection of the miscoverage rate while preserving finite-sample validity at the theoretical level (vovk2025conformal, gauthier2025values). Our algorithm synthesizes both advances by leveraging generative models to enhance decision-risk estimation accuracy and subsequently employing generalized inverse CP to assess the decision risk. However, we emphasize that our problem setting is fundamentally different and cannot be addressed by direct application of these methods. Substantive reformulation is required to adapt their underlying ideas to the decision-auditing context.
@@
Our work also connects to inverse optimization (IO), which seeks to infer the unknown parameters of an optimization problem from observed solutions (ahuja2001inverse, chan2025inverse, aswani2018inverse). Recent work on Conformal-IO combines IO with CP to prescribe robust decisions by first estimating decision-makers’ preference parameters from observed choices and then applies robust forward optimization using CP to construct an uncertainty set over these parameters (lin2024conformal, chan2024conformal). While we similarly integrate IO and CP, our learning setting fundamentally differs. We assume access to samples of the unknown parameters rather than the actual decisions. This leads us to employ inverse CP to assess decision reliability (singh2024distribution) rather than forward CP to prescribe decisions. Our focus on evaluation rather than prescription distinguishes our approach from Conformal-IO’s goal of generating policy recommendations. While efficient IO algorithms exist for structured problems like linear, integer, and convex programs (tavasliouglu2018structure, schaefer2009inverse, iyengar2005inverse), the lack of a unified framework for general IO leads to the adoption of structured formulations for tractability (chan2024conformal). This motivates our focus on linear models as part of our theoretical analysis, allowing us to derive an efficient closed-form solution for the proposed implementation algorithm.
@@
Finally, our research complements a growing line of work on human-algorithm collaboration in decision-making under uncertainty (kiyani2025decision, cortes2024decision, hullman2025conformal, cresswellconformal2). Unlike prior work that prescribes decisions, we leverage CP techniques to audit and assess the risk of candidate decisions, supporting human judgment through evaluation rather than recommendation. Additionally, our work contributes to the growing literature on the use of generative models to optimize in hybrid systems where human experts receive algorithmic advice (grand2024best, ibrahim2021eliciting, orfanoudaki2022algorithm). By producing a diverse set of plausible decisions, generative algorithms can act as advisory tools, thus promoting exploration and helping users develop more effective decision strategies (ajay2022conditional, li2024balancing).
@@ 3 Problem Setup
In particular, setting $\epsilon=0$ recovers the exact optimal decision rule. Throughout the manuscript, we suppress the subscript $\epsilon$ when no ambiguity arises, writing $\pi(y)$ for $\pi_{\epsilon}(y)$.
@@
The quantity $1-\alpha(z)$ serves as a conservative and distribution-free optimality certificate, providing a guaranteed confidence level that the prescribed decision $z$ is (near-)optimal under the true, unknown distribution $\mathcal{P}_{Y}$.
@@ 3.1 Motivating Examples
In this example, the human decision-maker acts based on an implicit perception of local weather patterns $Y$. The chosen decision $z$ thus reflects the decision-maker’s subjective belief about the likelihood of rain—carrying an umbrella if rain is perceived as more probable, and abstaining otherwise. Our goal is to assess whether this private knowledge aligns with the true local weather distribution by evaluating, through empirical data, how frequently the prescribed decision coincides with the optimal decision under the observed realizations of $Y$.
@@ 4 The CREDO Framework
We now present our method, which adapts the distribution-free uncertainty quantification framework of CP (angelopoulos2021gentle) to estimate decision risk through calibrated inverse feasible regions.
@@ 4.1 Preliminary: Conformal Prediction
CP (papadopoulos2002inductive) is a model-agnostic framework for uncertainty quantification that delivers finite-sample valid prediction regions under the exchangeability assumption: {assumption}[Exchangeability] The calibration data $\{(X_{i},Y_{i})\}_{i=1}^{n}$ and the test point $(X,Y)$ are exchangeable. That is, the joint distribution of $(X_{1},Y_{1}),\dots,(X_{n},Y_{n}),(X,Y)$ is invariant under any permutation of these $(n+1)$ pairs. Let the calibration dataset be $\mathcal{D}=\{(X_{i},Y_{i})\}_{i=1}^{n}$, drawn from an arbitrary joint distribution over $\mathcal{X}\times\mathcal{Y}$. For a user-specified miscoverage rate $\alpha$, the objective is to construct a prediction set $C(\cdot;\alpha)$ such that
@@
Overall architecture of the CREDO framework. The two main steps are: ($i$) Map the given candidate decision $z$ (e.g., $(0,0)$) to its inverse feasible region $\pi^{-1}(z)$; ($ii$) Assess the risk by constructing prediction sets within $\pi^{-1}(z)$ using generative conformal prediction, and extracting their miscoverage levels.
@@ 4.2 Conformalized Decision Risk Assessment
We now develop our framework for estimating the decision risk $\alpha(z)$ defined in Section 3. We accomplish this by reformulating the optimality probability in (2) as an inverse CP problem, where we seek to quantify how often uncertain parameters fall within regions that render a decision optimal. The key idea is to project the decision’s optimality condition onto an inverse feasible region (chan2025inverse, tavasliouglu2018structure) in the space of $\mathcal{Y}$, and to construct inner geometric approximations of this region using multiple generated conformal balls (shafer2008tutorial, wang2023probabilistic). The resulting average miscoverage rate across balls then serves as a data-driven lower-bound of decision risk. We detail this procedure in two steps, illustrated in Section 4.1.
@@ 4.2.1 Reformulation with Inverse Feasible Region
This definition generalizes its classical notion defined by chan2025inverse, providing a relaxed definition that accommodates near-optimal decisions. In parallel with our notation for $\pi$, we use $\pi^{-1}$ to denote $\pi_{\epsilon}^{-1}$ except where the distinction matters.
@@ 4.2.2 Risk Estimation via Generative Conformal Prediction
$\mathbb{P}\left\{Y\in\pi^{-1}(z)\right\}\stackrel{{\scriptstyle(i)}}{{\geq}}\mathbb{P}\left\{Y\in\mathcal{C}(X;\alpha)\right\}\stackrel{{\scriptstyle(ii)}}{{\geq}}1-\alpha.$ (6)

### v3

@@
In many operational settings, decision-makers must commit to actions before uncertainty resolves, but existing optimization tools rarely quantify how consistently a chosen decision remains optimal across plausible scenarios. This paper introduces CREDO—Conformalized Risk Estimation for Decision Optimization, a distribution-free framework that quantifies the probability that a prescribed decision remains (near-)optimal across realizations of uncertainty. CREDO reformulates decision risk through the inverse feasible region—the set of outcomes under which a decision is optimal—and estimates its probability using inner approximations constructed from conformal prediction balls generated by a conditional generative model. This approach yields finite-sample, distribution-free lower bounds on the probability of decision optimality. The framework is model-agnostic and broadly applicable across a wide range of optimization problems. Extensive numerical experiments demonstrate that CREDO provides accurate, efficient, and reliable evaluations of decision optimality across various optimization settings.
@@ 1 Introduction
Our proposed framework (CREDO) compared against conventional predict-then-optimize (PTO). For a given decision $z$ in feasible region $\mathcal{Z}$, CREDO first finds the inverse feasible region $\pi^{-1}(z)$ within scenario space $\mathcal{Y}$ and then estimates the associated decision risk by quantifying the probability that the realized $y$ lies within this region using conformal prediction (CP). In contrast, the conventional predict-then-optimize takes an estimated scenario $\hat{y}$ as input to policy $\pi$ and outputs a single prescribed decision without a reliability assessment.
To operationalize this perspective, we introduce CREDO—Conformalized Risk Estimation for Decision Optimization, a framework that quantifies, for any candidate decision, a distribution-free lower bound on the probability that it remains (near-)optimal. The proposed approach integrates inverse optimization (chan2025inverse) with distribution-free uncertainty quantification via conformal prediction (CP) (angelopoulos2021gentle) (see Figure 1 for an illustration). The core challenge is to evaluate the likelihood that an uncertain scenario falls within the decision’s inverse feasible region, namely the set of scenarios under which that decision remains optimal or near-optimal, and directly characterizing this region is generally intractable due to its implicit and complex geometry. CREDO circumvents this challenge by constructing inner approximations of the inverse feasible region using multiple CP balls generated from a conditional generative model. Each ball constitutes a valid prediction region for the outcome with guaranteed coverage. By shrinking the radius so that the ball lies entirely within the inverse feasible region, we obtain a certified lower bound on the decision’s optimality probability given by the corresponding coverage level. Averaging these calibrated bounds yields a distribution-free, finite-sample valid estimator of decision risk. CREDO is compatible with modern generative forecasting models, supports multiple conformal radius constructions, and applies broadly to convex decision problems. We establish theoretical validity for the estimator, develop efficient computational procedures, and demonstrate through extensive numerical experiments that CREDO provides accurate, interpretable, and reliable assessments of decision reliability across various optimization settings. The key contributions of our work are:
  • Problem Formulation: We formulate the decision risk assessment as estimating the probability that a candidate decision remains (near-)optimal under uncertainty. By expressing this probability through inverse feasible regions, we convert an intractable evaluation problem into a tractable probability estimation task.
  • CREDO Framework: We introduce CREDO, a distribution-free framework based on CP that delivers valid and finite-sample lower bounds on decision optimality, and establish marginal conservativeness, asymptotic consistency, and accuracy guarantees.
@@
  • Empirical Validation: Through controlled experiments featuring various optimization paradigms such as linear, quadratic, and second-order conic programs, as well as a real-world power grid planning application, we verify CREDO’s theoretical guarantees and demonstrate its practical advantages. Our method achieves 100% empirical validity under conservative radius choices, superior true positive rates (up to 78.75% improvement over baselines), consistently selects decisions with higher empirical confidence rankings than existing prescriptive methods, and is superior compared to various ablation variants.
The remainder of this paper is organized as follows. Section 2 reviews the related literature and positions our contributions. Section 3 formalizes the decision risk assessment problem and provides motivating examples. In Section 4, we present the CREDO methodology, combining inverse optimization with generative CP. Section 5 establishes theoretical guarantees including validity, consistency, and true positive rate analysis. Section 6 details the proposed computational implementation of CREDO, while Section 7 demonstrates its effectiveness through three numerical experiments. Section 8 concludes with discussions of implications and future work. Technical proofs and additional experiment details are included in the Appendix.
@@ 2 Related Work
Our work draws from and contributes to four distinct research streams: ($i$) decision-making under uncertainty in operations research, ($ii$) distribution-free uncertainty quantification through CP, (iiiiii) inverse optimization for understanding decision structures, and ($iv$) integration of CP and human-in-the-loop in robust decision making. In what follows, we position the CREDO approach relative to these lines of research.
@@
Classical approaches to optimization under uncertainty can be categorized by their treatment of unknown parameters. Stochastic optimization seeks decisions that minimize the expected value of the objective function across probabilistic scenarios, typically implemented through sample average approximation over historical data (shapiro2021lectures, kleywegt2002sample, lan2025bias). The predict-then-optimize (PTO) paradigm offers an alternative by first estimating unknown parameters via predictive models, then solving the resulting deterministic optimization problem (bertsimas2020predictive, elmachtoub2022smart). Acknowledging that probability distributions themselves may be misspecified, robust optimization (RO) is employed to hedge against worst-case realizations within specified uncertainty sets (bertsimas2006robust, ben2002robust). Distributionally robust optimization (DRO) generalizes this concept further, considering worst-case probability distributions rather than point realizations, thereby balancing conservativeness with statistical plausibility (delage2010distributionally, rahimian2022frameworks). Recognizing that prediction errors can compound when models are trained separately from their downstream use, decision-focused learning (DFL) trains predictive models end-to-end by differentiating through the optimization layer, directly minimizing downstream decision costs rather than prediction errors (amos2017optnet, mandi2024decision). While these prescriptive paradigms effectively recommend decisions, our work adopts a parallel goal: rather than prescribing a decision, we focus on auditing decisions by providing a quantitative, data-driven assessment of the risk associated with taking a given decision. This perspective complements existing decision-prescription methods and even human judgment by adding an additional layer of transparency and accountability, thereby supporting more reliable and defensible decision-making.
@@
Methodologically, our work builds upon the conformal prediction (CP) literature, a distribution-free approach for uncertainty quantification through calibrated prediction sets (shafer2008tutorial, vovk2005algorithmic). Multiple strands of recent research are relevant to our work. First, generative conformal methods draw multiple samples from generative models to construct both the calibration sets and the resulting CP regions, thus significantly improving efficiency (i.e., tightness of the prediction sets) especially when the data distribution is highly complex or dispersed (zheng2024optimizing, zhou2024hierarchical). Second, inverse CP estimates miscoverage rates for fixed prediction sets by identifying the smallest miscoverage level at which conformal regions are contained within the target set (prinster2023jaws, singh2024distribution). Advances in $e$-value CP strengthen this approach by enabling post-hoc selection of the miscoverage rate while preserving finite-sample validity at the theoretical level (vovk2025conformal, gauthier2025values). Our algorithm synthesizes both advances by leveraging generative models to enhance decision-risk estimation accuracy and subsequently employing generalized inverse CP to assess the decision risk. However, we emphasize that our problem setting is fundamentally different and cannot be addressed by direct application of these methods. Substantive reformulation is required to adapt their underlying ideas to the decision-auditing context.
@@
Our work also connects to inverse optimization (IO), which seeks to infer the unknown parameters of an optimization problem from observed solutions (ahuja2001inverse, chan2025inverse, aswani2018inverse, zattoni2025learning, besbes2025contextual). Recent work on Conformal-IO (lin2024conformal, chan2024conformal) combines IO with CP to prescribe robust decisions by first estimating decision-makers’ preference parameters from observed choices and then applying robust forward optimization using CP to construct an uncertainty set over these parameters. While we similarly integrate IO and CP, our learning setting fundamentally differs. We assume access to samples of the unknown parameters rather than the actual decisions. This leads us to employ inverse CP to assess decision reliability (singh2024distribution) rather than forward CP to prescribe decisions. Our focus on evaluation rather than prescription distinguishes our approach from Conformal-IO’s goal of generating policy recommendations. Meanwhile, though efficient IO algorithms exist for structured problems like linear, integer, and convex programs (tavasliouglu2018structure, schaefer2009inverse, iyengar2005inverse), the lack of a unified framework for general IO leads to the adoption of structured formulations for tractability (chan2024conformal). This motivates our focus on linear models as part of our theoretical analysis, allowing us to derive an efficient closed-form solution for the proposed implementation algorithm.
@@
Finally, our research complements a growing line of work on the integration of CP to robust decision-making (kiyani2025decision, patel2024conformal, cortes2024decision, yeh2024end, hullman2025conformal, cresswellconformal2, zhao2025conformal). Unlike prior work that prescribes decisions, we leverage CP techniques to audit and assess the risk of candidate decisions, supporting decision-making through evaluation rather than recommendation. Additionally, our work contributes to the growing literature on human-algorithm collaboration through the use of generative model recommendations (grand2024best, ibrahim2021eliciting, orfanoudaki2022algorithm). By producing a diverse set of plausible decisions, generative algorithms can act as advisory tools, thus promoting exploration and helping users develop more effective decision strategies (ajay2022conditional, li2024balancing).
@@ 3 Problem Setup
Note that setting $\epsilon=0$ recovers the exact optimal decision rule. Throughout the manuscript, we suppress the subscript $\epsilon$ when no ambiguity arises, writing $\pi(y)$ for $\pi_{\epsilon}(y)$.
@@
The quantity $1-\alpha(z)$ serves as a conservative optimality certificate, providing a guaranteed confidence level that the prescribed decision $z$ is (near-)optimal under the true, unknown distribution $\mathcal{P}_{Y}$.
@@ 3.1 Motivating Examples
In this example, the decision-maker acts based on an implicit perception of local weather patterns $Y$. The chosen decision $z$ thus reflects the decision-maker’s subjective belief about the likelihood of rain—carrying an umbrella if rain is perceived as more probable, and abstaining otherwise. Our goal is to assess whether this private knowledge aligns with the true local weather distribution by evaluating, through empirical data, how frequently the prescribed decision coincides with the optimal decision under the observed realizations of $Y$.
@@ 4 The CREDO Framework
We now present our method, which adapts the distribution-free uncertainty quantification framework of CP (vovk2005algorithmic) to estimate decision risk through calibrated inverse feasible regions.
@@ 4.1 Preliminary: Conformal Prediction
CP is a model-agnostic framework for uncertainty quantification that delivers finite-sample valid prediction regions under the exchangeability assumption: {assumption}[Exchangeability] The calibration data $\{(X_{i},Y_{i})\}_{i=1}^{n}$ and the test point $(X,Y)$ are exchangeable. That is, the joint distribution of $(X_{1},Y_{1}),\dots,(X_{n},Y_{n}),(X,Y)$ is invariant under any permutation of these $(n+1)$ pairs. Let the calibration dataset be $\mathcal{D}=\{(X_{i},Y_{i})\}_{i=1}^{n}$, drawn from an arbitrary joint distribution over $\mathcal{X}\times\mathcal{Y}$. For a user-specified miscoverage rate $\alpha$, the objective is to construct a prediction set $C(\cdot;\alpha)$ such that
@@
Example of the CREDO procedure applied to a linear programming problem. There are two main steps: ($i$) Map the given candidate decision $z$ (e.g., $(0,0)$) to its inverse feasible region $\pi^{-1}(z)$; ($ii$) Assess the risk by constructing inner approximation of $\pi^{-1}(z)$ using generative conformal prediction, and averaging their miscoverage levels.
@@ 4.2 Conformalized Decision Risk Assessment
We now develop our framework for estimating the decision risk $\alpha(z)$ defined in Section 3. We accomplish this by reformulating the optimality probability in (2) as an inverse CP problem, where we seek to quantify how often uncertain parameters fall within regions that render a decision optimal. The key idea is to project the decision’s optimality condition onto an inverse feasible region (chan2025inverse, tavasliouglu2018structure) in the space of $\mathcal{Y}$, and to construct inner geometric approximations of this region using multiple generated conformal balls (shafer2008tutorial, wang2023probabilistic). The resulting average miscoverage rate across balls then serves as a data-driven lower-bound of decision risk. We detail this procedure in two steps, with an example illustrated in Section 4.1.
@@ 4.2.1 Reformulation with Inverse Feasible Region
This definition can be viewed as a generalization of its classical notion defined by chan2025inverse, providing a relaxed definition that accommodates near-optimal decisions. For notation simplicity, we use $\pi^{-1}$ to denote $\pi_{\epsilon}^{-1}$ when clear from context. The reformulation is formally stated as follows.
@@ 4.2.2 Risk Estimation via Generative Conformal Prediction
$\mathbb{P}\left\{Y\in\pi^{-1}(z)\right\}\stackrel{{\scriptstyle(i)}}{{\geq}}\mathbb{P}\left\{Y\in\mathcal{C}(X;\alpha)\right\}\stackrel{{\scriptstyle(ii)}}{{\geq}}1-\alpha.$